OpenStudy (annie96):

Anita and Joelle bowl together and their combined total score for one game was 425 points. Anita’s score was 40 less than twice Joelle’s. What were their scores? Write a system of equations to model the problem if x represents Joelle’s score, and y represents Anita’s score. (Points : 1) x + y = 425 y = 2x – 40 x + y = 425 y = 2x + 40 x – y = 425 y = 2x – 40 x + y = 40 y = 2x – 425

OpenStudy (anonymous):

The reason is, because the combined scores make 425, it's x+y=425 The second part is because 40 less that 2 times means 2 times-40, and that can be written: y=2x-40 So, the answer is A.

OpenStudy (anonymous):

Joelle's Score: 155 Anita's Score: 270 155*2=310 310-40=270 270+155=425